OpenStudy (anonymous):

Find the circumference of a circle radius a(i.e find the arc length of the circle of radius a) Hint: integration

OpenStudy (anonymous):

Let \(\tau = 2\pi\) \(s = r \theta\) \(\theta = \tau\) \(\Rightarrow s = r \tau\)